OpenStudy (anonymous):

Identify the resulting probability as theoretical or experimental. One-half of the people in a chess club are adults 21 years old or older. The other half are younger. What is the probability that a club member chosen at random will be under 21 years old? A. The experimental probability is 50%. B. The experimental probability is 75%. C. The theoretical probability is 75%. D. The theoretical probability is 50%.

OpenStudy (anonymous):

@Whitemonsterbunny17

OpenStudy (anonymous):

@Kristen17 @Kkutie7

OpenStudy (kkutie7):

what percent does one-half represent? hint \[\\frac{1}{2}] divide them and you get a decimal point... multiply it by 100 see what you get =)

OpenStudy (anonymous):

WaitWhat? LOL @Kkutie7

OpenStudy (kkutie7):

oh shoot *\[\frac{1}{2}\] because this is one half right

OpenStudy (anonymous):

I dont know haha @Kkutie7

OpenStudy (kkutie7):

one half is the same as saying \[\frac{1}{2}\] going from this knowledge you can divide them to get a decimal. you can use a calculator.

OpenStudy (anonymous):

Do I divide 1/2 with 100?

OpenStudy (kkutie7):

no just divide 1 by two right now ok... first step =)

OpenStudy (anonymous):

Soi did 1 divided by 2 and i got 0.5

OpenStudy (kkutie7):

ok good to get percent you multiply that by 100

OpenStudy (anonymous):

duh its 50 :)

OpenStudy (anonymous):

@Kkutie7

OpenStudy (kkutie7):

50% is half =) so if the class has one-half or 50% under 21 the odds are 50% that you will get someone under 21

OpenStudy (anonymous):

Okay so is the answer A?

OpenStudy (anonymous):

@Kkutie7

OpenStudy (kkutie7):

yupp =)

OpenStudy (anonymous):

What is the diference between A and D though?

OpenStudy (kkutie7):

A means you could test it and that would be the answer, theoretically means in theory... or by thought. I would go with that one to be honest.

OpenStudy (anonymous):

K TYSMMM